One way the U.S. Environmental Protection Agency (EPA) tests for chloride contaminants in water is by titrating a sample of silver nitrate solution. Any chloride anions in solution will combine with the silver cations to produce bright white silver chloride precipitate. Suppose an EPA chemist tests a sample of groundwater known to be contaminated with nickel(II) chloride, which would react with silver nitrate solution like this: (aq) (aq) (s) (aq) The chemist adds M silver nitrate solution to the sample until silver chloride stops forming. He then washes, dries, and weighs the precipitate. He finds he has collected of silver chloride. Calculate the concentration of nickel(II) chloride contaminant in the original groundwater sample. Round your answer to significant digits.
Answer #1

Here, no values are given.

So, let's consider you have a 500 mL solution of 0.01 M NiCl2 and you treated and you treat with AgNO3 solution, according to the following equation.

NiCl2(aq) + 2AgNO3(aq) ----> Ni(NO3)2(aq) + 2AgCl(s)

Let's say you obtained 1.435 g of AgCl solid.

The molar mass of AgCl = 143.5 g/mol

i.e. The moles of AgCl = 1.435 g/(143.5 g/mol) = 0.01 mol

Hence, the moles of chloride (Cl-) = 0.01 mol

Now, the moles of NiCl2 = 0.01/2 = 0.005 mol

Therefore, the concentration of NiCl2 = 0.005 mol/(500/1000) L = 0.01 mol/L or 0.01 M